OpenStudy (anonymous):

c-2/12=3/9 what is c.

hero (hero):

cross multiply

hero (hero):

9(c-2) = 36 9c - 18 = 36 9c = 54 c = 54/9 c = 6

OpenStudy (anonymous):

c=6

OpenStudy (anonymous):

if that were a fraction 1/6?

OpenStudy (anonymous):

c=18